OpenStudy (anonymous):

In the system shown below, what is the sum of the x-coordinates of all solutions? 3y^2-5x^2=7 3y^2-x^2=23

OpenStudy (anonymous):

3y^2 - 5x^2 = 7 3y^2 = 7 + 5x^2 y = √[(7 + 5x^2)/3] 3y^2 - x^2 = 23 3√[(7 + 5x^2)/3]^2 - x^2 = 23 (7 + 5x^2) - x^2 = 23 4x^2 = 16 x^2 = 4 x = -2 [Not +2 because in the second quadrant, x is negative] y = √[(7 + 5x^2)/3] y = √[(7 + 5(-2)^2)/3] y = √[(7 + 20)/3] y = √9 y = +3 [Not -3 because in the 2nd quadrant y is positive] (-2,3) you asked this ques... once again!! but hope this helps :-)

OpenStudy (callisto):

Well, neglect what the above person said.. I don't really know how come the term '' second quadrant'' appears :S To do this, just solve for x and add the x value together 3y^2-5x^2=7 -(1) 3y^2-x^2=23 => 3y^2 = 23 +x^2 -(2) Put (2) into (1) 23 + x^2 -5x^2=7 -4x^2 = -16 x^2 = 4 Can you solve it from here and get your answer?

OpenStudy (anonymous):

can you help me out a little more please

OpenStudy (callisto):

x = 2 or x=-2 Can you do it?

OpenStudy (anonymous):

-2?

OpenStudy (callisto):

sum of the x-coordinates of all solutions For that i think you should add all the x values together?

OpenStudy (anonymous):

ah I'm sorry but can you just be straight forward with me I'm tired and haven't gone to bed yet :/

OpenStudy (callisto):

First, I'm not sure if it is like that x-coordinates of all solutions = sum of x values = 2+ (-2) =0 I don't really know it :S

OpenStudy (anonymous):

sum of the x - coordinate is zero

OpenStudy (anonymous):

may you please rewrite again the system of eqations